# GDSG: Graph Diffusion-based Solution Generation for Optimization Problems in MEC Networks

### v1

@@ IV The Proposed GDSG Algorithm
With the solution distribution parametrization provided, the core theoretical concept in Sec. III is to establish a model capable of generating $\theta^{\prime}\sim p(\Theta_{x})$ for a given $x$. Generative models aim to model the joint distribution of inputs and outputs, $P(\Theta_{x},x)$, in contrast to discriminative models, which learn the conditional probability distribution $P(\Theta_{x}|x)$. Consequently, the generative model uses the implicitly learned $P(\Theta_{x}|x)$ as a guide to generate $\theta^{\prime}\sim p(\Theta_{x})$ in conditional generation [33], thereby enhancing the likelihood of producing high-quality outputs. To address the graph optimization problem outlined in Sec. II and implement the approach outlined in Sec. III, we propose a graph diffusion generative model for solution generation.
@@ IV-A Graph Diffusion for Solution Generation
From the perspective of variational inference, the general diffusion framework consists of a forward denoising process and a learnable reverse denoising Markov process [32, 50, 51].
As described in Sec. II, for an input $\mathcal{G}(\mathcal{V},\mathcal{E})$, the optimal solution to the multi-server multi-user computation offloading problem can be expressed as $\{\mathcal{D}^{*},\mathcal{A}^{*}\}$. Next, we introduce the diffusion generation of $\mathcal{D}^{*}$ and $\mathcal{A}^{*}$ in the following two aspects: discrete solutions and continuous solutions generation, and the output solutions are represented by $\mathrm{y}$ and $y$ for discrete and continuous solution respectively.
@@ IV-A1 Discrete Solution Diffusion
For discrete solution $\mathrm{y}\in\mathcal{Y}$, it is parametrized as a one-hot vector that $\mathrm{y}\in\{0,1\}^{L\times 2}$. Given an instance $(\mathcal{G},\mathrm{y})$, diffusion takes $\mathrm{y}$ as the initial solution $\mathrm{y}_{0}$ to perform the denoising process, also known as the diffusion process, gradually introducing noise to obtain a series of variables $\mathrm{y}_{1:T}=\mathrm{y}_{1},\ldots,\mathrm{y}_{T}$ with $T$ as the diffusion steps. The specific denoising process can be formalized as $q(\mathrm{y}_{1:T}|\mathrm{y}_{0})=\prod^{T}_{t=1}q(\mathrm{y}_{t}|\mathrm{y}_{t-1})$. The model learns the noise distribution from this process, enabling it to denoise from a given initial noise distribution to a target distribution. The denoising process actually starts from $\mathrm{y}_{T}$ and performs adjacent step denoising from $\mathrm{y}_{t}$ to $\mathrm{y}_{t-1}$ with $\mathcal{G}$ as the condition. The denoising process can be expressed as $p_{\Phi}(\mathrm{y}_{0:T}|\mathcal{G})=p(\mathrm{y}_{T})\prod^{T}_{t=1}p_{\Phi}(\mathrm{y}_{t-1}|\mathrm{y}_{t},\mathcal{G})$, where $\Phi$ represents the model parameters. The ideal result of diffusion model training is $p_{\Phi}(\mathrm{y}|\mathcal{G})=q(\mathrm{y}|\mathcal{G})$, and to fit the original data distribution, the variational upper bound of the negative log-likelihood is generally used as the loss objective
$\mathcal{L}=\mathbb{E}_{q}\big{[}-\mathrm{log}p_{\Phi}(\mathrm{y}_{0}|\mathrm{y}_{1},\mathcal{G})$
$+\prod_{t>1}D_{KL}[q(\mathrm{y}_{t-1}|\mathrm{y}_{t},\mathrm{y}_{0})\ ||\ p_{\Phi}(\mathrm{y}_{t-1}|\mathrm{y}_{t},\mathcal{G})]\big{]}+C,$ (14)
@@
In the forward noising process, $\mathrm{y}_{t}$ is multiplied by $\mathbf{Q}_{t}$ to obtain $\mathrm{y}_{t+1}$, where $\mathbf{Q}_{t}=\begin{bmatrix}(1-\beta_{t})&\beta_{t}\\ \beta_{t}&(1-\beta_{t})\end{bmatrix}$ is the transition probability matrix [35, 52]. The operation $\mathrm{y}\mathbf{Q}$ converts and adds noise to the one-hot vector of each element $\mathrm{y}_{i}\in\{0,1\}^{2}$, with $\beta_{t}$ controlling the noise scale as the corruption ratio. Specifically, $\prod^{T}_{t=1}(1-\beta_{t})\approx 0$ ensures that $\mathrm{y}_{T}\sim\mathrm{Uniform}(\cdot)$. Let $\overline{\mathbf{Q}}_{t}=\mathbf{Q}_{1}\mathbf{Q}_{2}\ldots\mathbf{Q}_{t}$, the single-step and t-step marginals of the noising process can be formulated as
$q(\mathrm{y}_{t}|\mathrm{y}_{t-1})$ $=\mathrm{Cat}(\mathrm{y}_{t};\mathbf{p}=\mathrm{y}_{t-1}\mathbf{Q})\ \ \ \mathrm{and}$
$q(\mathrm{y}_{t}|\mathrm{y}_{0})$ $=\mathrm{Cat}(\mathrm{y}_{t};\mathbf{p}=\mathrm{y}_{0}\overline{\mathbf{Q}}_{t}),$ (15)
@@
$q(\mathrm{y}_{t-1}|\mathrm{y}_{t},\mathrm{y}_{0})=\frac{q(\mathrm{y}_{t}|\mathrm{y}_{t-1},\mathrm{y}_{0})q(\mathrm{y}_{t-1}|\mathrm{y}_{0})}{q(\mathrm{y}_{t}|\mathrm{y}_{0})}$
$=\mathrm{Cat}(\mathrm{y}_{t-1};\mathbf{p}=\frac{\mathrm{y}_{t}\mathbf{Q}^{T}_{t}\odot\mathrm{y}_{0}\overline{\mathbf{Q}}_{t-1}}{\mathrm{y}_{0}\overline{\mathbf{Q}}_{t}\mathrm{y}^{T}_{t}}),$ (16)
@@
The model is trained to predict the distribution $p_{\Phi}(\tilde{\mathrm{y}}_{0}|\mathrm{y}_{t},\mathcal{G})$, allowing the denoising process to be obtained by substituting the $\mathrm{y}_{0}$ in Eq. (IV-A1) with the estimated $\tilde{\mathrm{y}}_{0}$
$p_{\Phi}(\mathrm{y}_{t-1}|\mathrm{y}_{t})\propto\sum_{\tilde{\mathrm{y}}_{0}}q(\mathrm{y}_{t-1}|\mathrm{y}_{t},\tilde{\mathrm{y}}_{0})p_{\Phi}(\tilde{\mathrm{y}}_{0}|\mathrm{y}_{t},\mathcal{G}).$ (17)
@@ IV-A2 Continuous Solution Diffusion
For the continuous solution $y\in[0,1]^{L}$, continuous diffusion models [32, 33, 34] can be directly applied without parametrization. Unlike previous studies [36, 37], which rescales discrete data in continuous diffusion for discrete data generation [53], our multi-task generation approach uses continuous diffusion for continuous solutions.
@@
$q(y_{t}|y_{t-1})$ $=\mathcal{N}(y_{t};\sqrt{1-\beta_{t}}y_{t-1},\beta_{t}\mathbf{I})$
$q(y_{t}|y_{0})$ $=\mathcal{N}(y_{t};\sqrt{\overline{(1-\beta_{t})}}y_{0},(1-\overline{(1-\beta_{t})})\mathbf{I}),$ (18)
where $\overline{(1-\beta_{t})}=\prod^{t}_{t=1}(1-\beta_{t})$, and $\prod^{T}_{t=1}(1-\beta_{t})\approx 0$ ensures that $\mathrm{y}_{T}\sim\mathcal{N}(\cdot)$. Similar to Eq. (IV-A1), the posterior can be calculated as
$q(y_{t-1}|y_{t},y_{0})=\frac{q(y_{t}|y_{t-1},y_{0})q(y_{t-1}|y_{0})}{q(y_{t}|y_{0})}.$ (19)
The model is trained to predict the Gaussian noise $\tilde{\mathbf{\epsilon}}_{t}=f_{\Phi}(y_{t},t,\mathcal{G})$. Given that $\tilde{\mathbf{\epsilon}}_{t}=(y_{t}-\sqrt{\overline{(1-\beta_{t})}}y_{0})/\sqrt{1-\overline{(1-\beta_{t})}}$, the denoising process can be obtained by substituting the $y_{0}$ in Eq. (19) with an estimation
$p_{\Phi}(y_{t-1}|y_{t})\propto q(y_{t-1}|y_{t},\frac{y_{t}-\sqrt{1-\overline{(1-\beta_{t})}}f_{\Phi}(y_{t},t,\mathcal{G})}{\sqrt{\overline{(1-\beta_{t})}}}).$ (20)
@@ IV-B GNN-based Diffusion Model
However, we noticed that the existing works [36, 37] using convolutional GNN only considered fully connected input graphs for the TSP and MIS problems, whereas most graph optimization problems are not fully connected. In dense adjacency matrix and sparse edge list graph data structures, some edges may not exist, which we refer to as padding edges. The feature encoding of these padding edges can have unpredictable (and often adverse) effects on the iteration of the GNN in the latent space. This issue is not well addressed by previous works.
We investigated this problem during our experiments, as detailed in Sec. V-D. Specifically, the feature embedding of padding edges is non-zero. The input graphs of different sizes may have varying proportions of padding edges, and their influence is difficult to generalize or eliminate through simple linear or nonlinear gating. Changing the aggregation method (e.g., sum, mean, max) offers little help. Moreover, we considered setting the original features of padding edges to zero and setting the bias of the edge feature embedding layer to zero simultaneously. However, this approach could adversely affect the embedding of normal edges and might still not completely distinguish padding edges from normal edges in some cases.
To address this, we design a gating mechanism that can mitigate the effects of padding edges, inspired by the spirits that the generalization performance of GNNs benefits from explicit redesign [54, 56]. We add a one-dimensional 0-1 variable to the edge feature to indicate whether an edge is a padding edge. This variable multiplies the original gating result, completely eliminating the influence of meaningless hidden features from padding edges.
At this point, our GNN model has been extended to multi-task capabilities and improved in generalization for both fully connected and non-fully connected graphs. Additionally, as an anisotropic GNN handling directionally dependent features, our model is not limited to the directed graph problem in Sec. II and can be directly applied to undirected graphs. GNNs inherently offer versatility, allowing a single trained model to handle graphs of different scales for the same problem. Moreover, the settings of node and edge features can be flexibly adjusted to support various MEC tasks.
@@ IV-C Multi-Task Generation and Denoising Schedule
Our model performs multi-task generation, denoising both the target discrete solution and the continuous solution. In multi-task learning, negative competition between tasks can lead to poor convergence. However, we find that the discrete and continuous diffusion tasks largely satisfy the metric of inter-task gradient direction orthogonality, a key goal in specialized multi-task learning research [57, 58, 59]. Our experiments reveal that the gradient vectors calculated on the same layer parameters for the two tasks are almost orthogonal. Details are provided in Sec. V-C. A comparison using the same neural network structure for discriminative training confirms that this orthogonality is not inherent to the network architecture. We believe this characteristic is unique to the training loss objectives of discrete and continuous diffusion, a phenomenon not yet reflected in existing multi-task learning research on diffusion models [60, 61, 62].
Conventionally, we employ DDIM to accelerate the denoising inference of the diffusion model [63]. DDIM rewrites the denoising formula, leveraging the independence of marginal probabilities between two adjacent steps in the original process. This approach significantly reduces the number of inference steps while maintaining denoising performance. For discrete diffusion, we use a similar method [35].
@@ V Experiments
To verify our proposed methods and models, we build instances of the multi-server multi-user computation offloading problem at various scales. We first provide targeted simulation validation for the approach outlined in Sec. III, then we conduct a gradient analysis for multi-task orthogonality and compare GDSG with a discriminative model of the same neural network structure. Finally, we compare GDSG with multiple baselines in terms of cost metrics.

### v3

@@ IV The Proposed GDSG Algorithm
With the solution distribution parametrization provided, the core theoretical concept in Sec. III is to establish a model capable of generating $\boldsymbol{\theta}^{\prime}\sim p(\boldsymbol{\Theta}_{\mathbf{x}})$ for a given $\mathbf{x}$. Generative models aim to model the joint distribution of inputs and outputs, $P(\boldsymbol{\Theta}_{\mathbf{x}},\mathbf{x})$, in contrast to discriminative models, which learn the conditional probability distribution $P(\boldsymbol{\Theta}_{\mathbf{x}}|\mathbf{x})$. Consequently, the generative model uses the implicitly learned $P(\boldsymbol{\Theta}_{\mathbf{x}}|\mathbf{x})$ as a guide to generate $\boldsymbol{\theta}^{\prime}\sim p(\boldsymbol{\Theta}_{\mathbf{x}})$ in conditional generation [40], thereby enhancing the likelihood of producing high-quality outputs. To address the graph optimization problem outlined in Sec. II and implement the approach outlined in Sec. III, we propose a graph diffusion generative model for solution generation. The concepts and analysis in Section III can be applied to various generative models. However, due to the diffusion model’s capability to accommodate different data types and its versatility across various model scales, we focus on GDM in this work.
@@ IV-A Graph Diffusion for Solution Generation
From the perspective of variational inference, the general diffusion framework consists of a forward-noising process and a learnable reverse denoising Markov process [39, 60, 61].
As described in Sec. II, for an input $\mathcal{G}(\mathcal{V},\mathcal{E})$, the optimal solution to the multi-server multi-user computation offloading problem can be expressed as $\{\mathcal{D}^{*},\mathcal{A}^{*}\}$. Next, we introduce the diffusion generation of $\mathcal{D}^{*}$ and $\mathcal{A}^{*}$ in the following two aspects: discrete solutions and continuous solutions generation.
@@ IV-A1 Discrete Solution Diffusion
For discrete solution $\mathbf{y}\in\mathcal{Y}$, it is parametrized as a one-hot vector that $\mathbf{y}\in\{0,1\}^{L\times 2}$. Given an instance $(\mathcal{G},\mathbf{y})$, diffusion takes $\mathbf{y}$ as the initial solution $\mathbf{y}_{0}$ to perform the noising process, also known as the diffusion process, gradually introducing noise to obtain a series of variables $\mathbf{y}_{1:T}=\mathbf{y}_{1},\ldots,\mathbf{y}_{T}$ with $T$ as the diffusion steps. The specific noising process can be formalized as $q(\mathbf{y}_{1:T}|\mathbf{y}_{0})=\prod^{T}_{t=1}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$. The model learns the noise distribution from this process, enabling it to denoise from a given initial noise distribution to a target distribution. The denoising process actually starts from $\mathbf{y}_{T}$ and performs adjacent step denoising from $\mathbf{y}_{t}$ to $\mathbf{y}_{t-1}$ with $\mathcal{G}$ as the condition. The denoising process can be expressed as $p_{\Phi}(\mathbf{y}_{0:T}|\mathcal{G})=p(\mathbf{y}_{T})\prod^{T}_{t=1}p_{\Phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathcal{G})$, where $\Phi$ represents the model parameters. The ideal result of diffusion model training is $p_{\Phi}(\mathbf{y}|\mathcal{G})=q(\mathbf{y}|\mathcal{G})$, and to fit the original data distribution, the variational upper bound of the negative log-likelihood is generally used as the loss objective
$\mathcal{L}=\mathbb{E}_{q}\big{[}-\mathrm{log}p_{\Phi}(\mathbf{y}_{0}|\mathbf{y}_{1},\mathcal{G})$
$+\prod_{t>1}D_{KL}[q(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0})\ ||\ p_{\Phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathcal{G})]\big{]}+C,$ (14)
@@
In the forward noising process, $\mathbf{y}_{t}$ is multiplied by $\mathbf{Q}_{t}$ to obtain $\mathbf{y}_{t+1}$, where $\mathbf{Q}_{t}=\begin{bmatrix}(1-\beta_{t})&\beta_{t}\\ \beta_{t}&(1-\beta_{t})\end{bmatrix}$ is the transition probability matrix [42, 62]. The operation $\mathbf{y}\mathbf{Q}$ converts and adds noise to the one-hot vector of each element $y_{i}\in\{0,1\}^{2}$, with $\beta_{t}$ controlling the noise scale as the corruption ratio. Specifically, $\prod^{T}_{t=1}(1-\beta_{t})\approx 0$ ensures that $\mathbf{y}_{T}\sim\mathrm{Uniform}(\cdot)$. Let $\overline{\mathbf{Q}}_{t}=\mathbf{Q}_{1}\mathbf{Q}_{2}\ldots\mathbf{Q}_{t}$, the single-step and $t$-step marginals of the noising process are formulated as
$q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$ $=\mathrm{Cat}(\mathbf{y}_{t};\mathbf{p}=\mathbf{y}_{t-1}\mathbf{Q})\ \ \ \mathrm{and}$
$q(\mathbf{y}_{t}|\mathbf{y}_{0})$ $=\mathrm{Cat}(\mathbf{y}_{t};\mathbf{p}=\mathbf{y}_{0}\overline{\mathbf{Q}}_{t}),$ (15)
@@
$q(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0})=\frac{q(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{y}_{0})q(\mathbf{y}_{t-1}|\mathbf{y}_{0})}{q(\mathbf{y}_{t}|\mathbf{y}_{0})}$
$=\mathrm{Cat}(\mathbf{y}_{t-1};\mathbf{p}=\frac{\mathbf{y}_{t}\mathbf{Q}^{T}_{t}\odot\mathbf{y}_{0}\overline{\mathbf{Q}}_{t-1}}{\mathbf{y}_{0}\overline{\mathbf{Q}}_{t}\mathbf{y}^{T}_{t}}),$ (16)
@@
The model is trained to predict the distribution $p_{\Phi}(\tilde{\mathbf{y}}_{0}|\mathbf{y}_{t},\mathcal{G})$, allowing the denoising process to be obtained by substituting the $\mathbf{y}_{0}$ in Eq. (IV-A1) with the estimated $\tilde{\mathbf{y}}_{0}$
$p_{\Phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t})\propto\sum_{\tilde{\mathbf{y}}_{0}}q(\mathbf{y}_{t-1}|\mathbf{y}_{t},\tilde{\mathbf{y}}_{0})p_{\Phi}(\tilde{\mathbf{y}}_{0}|\mathbf{y}_{t},\mathcal{G}).$ (17)
@@ IV-A2 Continuous Solution Diffusion
For the continuous solution $\mathbf{y}\in[0,1]^{L}$, continuous diffusion models [39, 40, 41] can be directly applied without parametrization. Unlike previous studies [43, 44] rescaling discrete data in continuous diffusion for discrete data generation [63], our multi-task generation approach uses continuous diffusion for continuous solutions.
@@
$q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$ $=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-\beta_{t}}\mathbf{y}_{t-1},\beta_{t}\mathbf{I})$
$q(\mathbf{y}_{t}|\mathbf{y}_{0})$ $=\mathcal{N}(\mathbf{y}_{t};\sqrt{\overline{(1-\beta_{t})}}\mathbf{y}_{0},(1-\overline{(1-\beta_{t})})\mathbf{I}),$ (18)
where $\overline{(1-\beta_{t})}=\prod^{t}_{t=1}(1-\beta_{t})$, and $\prod^{T}_{t=1}(1-\beta_{t})\approx 0$ ensures that $\mathbf{y}_{T}\sim\mathcal{N}(\cdot)$. Similar to Eq. (IV-A1), the posterior can be calculated as
$q(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0})=\frac{q(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{y}_{0})q(\mathbf{y}_{t-1}|\mathbf{y}_{0})}{q(\mathbf{y}_{t}|\mathbf{y}_{0})}.$ (19)
The model is trained to predict the Gaussian noise $\tilde{\mathbf{\boldsymbol{\epsilon}}}_{t}=f_{\Phi}(\mathbf{y}_{t},t,\mathcal{G})$. Given that $\tilde{\mathbf{\boldsymbol{\epsilon}}}_{t}=(\mathbf{y}_{t}-\sqrt{\overline{(1-\beta_{t})}}\mathbf{y}_{0})/\sqrt{1-\overline{(1-\beta_{t})}}$, the denoising process can be obtained by substituting the $\mathbf{y}_{0}$ in Eq. (19) with an estimation
$p_{\Phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t})\propto q(\mathbf{y}_{t-1}|\mathbf{y}_{t},\frac{\mathbf{y}_{t}-\sqrt{1-\overline{(1-\beta_{t})}}f_{\Phi}(\mathbf{y}_{t},t,\mathcal{G})}{\sqrt{\overline{(1-\beta_{t})}}}).$ (20)
@@ IV-B GNN-based Diffusion Model
However, we notice that the existing works [43, 44] using convolutional GNN only consider fully connected input graphs for the TSP and MIS problems, whereas most graph optimization problems are not fully connected. In dense adjacency matrix and sparse edge list graph data structures, some edges may not exist, which we refer to as padding edges. The feature encoding of these padding edges can have unpredictable (and often adverse) effects on the iteration of the GNN in the latent space. This issue is not well addressed by previous works.
We investigate this problem during our experiments, as detailed in Sec. V-D. Specifically, the feature embedding of padding edges is non-zero. The input graphs of different sizes may have varying proportions of padding edges, and their influence is difficult to generalize or eliminate through simple linear or nonlinear gating. Changing the aggregation method (e.g., sum, mean, max) offers little help. Moreover, we considered setting the original features of padding edges to zero and setting the bias of the edge feature embedding layer to zero simultaneously. However, this approach could adversely affect the embedding of normal edges and might still not completely distinguish padding edges from normal edges in some cases.
To address this, we design a gating mechanism that can mitigate the effects of padding edges, inspired by the spirit that the generalization performance of GNNs benefits from explicit redesign [64, 66]. We add a one-dimensional 0-1 variable to the edge feature to indicate whether an edge is a padding edge. This variable multiplies the original gating result, eliminating the influence of meaningless hidden features from padding edges. Notably, the settings of node and edge features can be flexibly adjusted to support various MEC tasks.
The proposed GDSG method can be extended to multi-task capabilities and improve generalization for both fully connected and non-fully connected graphs. Additionally, as an anisotropic GNN handling directionally dependent features, our model is not limited to the directed graph problem and can be also applied to undirected graphs. GNNs inherently offer versatility, allowing a single trained model to handle graphs of different scales for the same problem.
@@ IV-C Multi-Task Generation and Denoising Schedule
Our model is designed for multi-task generation, specifically focusing on denoising both the target discrete solution and the continuous solution. In the area of multi-task learning [67, 68], it is recognized that negative competition between tasks can impede convergence. During training, the gradient directions of different task losses are often misaligned when they share the same parameters. Simply combining these non-orthogonal gradients may lead to suboptimal training results, sometimes causing worse performance on individual tasks compared to single-task training. Our approach, GDSG, incorporates both continuous and discrete diffusion, with the two tasks sharing all parameters except for the output heads. Therefore, we analyze this interaction carefully to mitigate potential degradation in training performance.
However, we discover that the discrete and continuous diffusion tasks primarily align with the principle of inter-task gradient direction orthogonality, which is a significant goal in specialized multi-task learning research [67, 68, 69]. We find that the gradient vectors calculated for these two tasks, using the same layer parameters, are nearly orthogonal. This finding is further elaborated in Section V-C. A comparative study using the same neural network architecture for discriminative training indicates that this orthogonality does not stem from the network architecture itself. We posit that this characteristic is intrinsic to the training loss objectives for discrete and continuous diffusion, representing a phenomenon not yet captured in the current multi-task learning research focused on diffusion models [70, 71, 72]. We observe that task orthogonality is not only an impressive side effect of our method, it also demonstrates the GDM’s great potential for scalability in tasks.
Additionally, we employ DDIM to accelerate the denoising inference of the diffusion model [73]. DDIM rewrites the denoising formula, leveraging the independence of marginal probabilities between two adjacent steps in the original process. This approach significantly reduces the number of inference steps while maintaining denoising performance.
@@ V Experiments
To verify our proposed methods and models, we establish instances of the MSCO problem at various scales. We first provide targeted simulation validation for the approach outlined in Sec. III, and then we conduct a gradient analysis for multi-task orthogonality and compare GDSG with a discriminative model of the same neural network structure. Finally, we compare GDSG with multiple baselines in terms of cost metrics.